Today’s activities: Met obs + GPS trial continues. Snow sampling. Despite several brief partial clearances conditions remain poor with visibility down to less than 10 meters for most of the day. Cats kept running round the clock with 2hrly checks. General caboose maintenance. All OK.

Intentions for tomorrow: Normal obs. Domestic scrubout. Conditions permitting we will start moving fuel scoots 5km north.

Long term intentions: Ice Train will start moving as 6 units – 4 x units of fuel scoots. 1 x unit Living Caboose. 1 x unit Science Caboose + ISO storage container. First move will be 5km. Second move will be 4 km which then puts us on the edge of a significant crevasse area. 5 days of White Mars sampling begins on Sun 22nd.